OpenStudy (anonymous):

Can anyone show me the way to solve this one without graphing calculator. 1/4 + sin(pi*x) = 4^(-x). Solve for x. I need help please

OpenStudy (anonymous):

This is a tricky question. What are your ideas?

OpenStudy (amistre64):

you can start by narrowing down the domain

OpenStudy (amistre64):

maybe redefine sin in terms of a polynomial?

OpenStudy (amistre64):

or isnt there an e version of sin?

OpenStudy (anonymous):

Ok sure. That might work and if it's the right approach it's beyond what I know how to do off the top of my head. One thing to do, even though you need to find the solution without a graphing calculator is to sketch both functions to see where they might meet. To get an idea of if there are any solutions or just one or many. Just get an idea as to what is going on.

OpenStudy (anonymous):

(check your definitions and identities from the course or the chapter ... there are many ways to rewrite functions)

OpenStudy (anonymous):

They don't have specific answer and all the equation may beyond my level.

OpenStudy (amistre64):

\[4^x(\frac{1}{4} + \sin(pi*x) = 4^{-x})\] \[4^{x-1} + 4^x\sin(pi*x) = 1\] is some ways to move things about

OpenStudy (anonymous):

here's a picture of the two function on top of each other. so clearly they meet at multiple points.

OpenStudy (anonymous):

\[\frac{ 1 }{ 4 } + \sin (pix) = 4^{-x}\]

OpenStudy (anonymous):

This is what the question about

OpenStudy (amistre64):

\[sin(ab)=sin(a)cos(b)+sin(b)cos(a)\] \[sin(pi~ x)=sin(pi)cos(x)+sin(x)cos(pi)\] \[sin(pi~ x)=-sin(x)\]

OpenStudy (amistre64):

4^-x is never zero ...

OpenStudy (anonymous):

if I have graphing calculator I can see the point where they are intersect. The main idea is I need help to solve for x without graphing calculator mrleiss

OpenStudy (anonymous):

I know 4^-x will never be zero

OpenStudy (amistre64):

this is a process, you determine what you can do with the knowledge you have at your disposal.

OpenStudy (anonymous):

Ok. The other idea is to plug in some values that are special to both exponential functions and trig (or sine in particular) and see if you get any that are the same or close. What values should you try?

OpenStudy (anonymous):

This is a free respond question and they don't give you any number to plug in

OpenStudy (amistre64):

that trig identity is off, sin (a+b) not ab :/

OpenStudy (anonymous):

I know, but what values should you always try - they are the easy ones!

OpenStudy (anonymous):

you can't tried any value if you don't know what you are doing

OpenStudy (anonymous):

it is definitely not sin(a+b) I swear

OpenStudy (anonymous):

You can, and should, think about trying 0, 1, -1 for just about any function or equation; especially if you don't know where to start.

OpenStudy (anonymous):

are you in calculus?

OpenStudy (anonymous):

It will help you, along with a sketch, to get a feel for what the question is about. You should be able to sketch y = 4^(-x) and y = sin(pi*x) without too much trouble. I put it in the graphing calculator because it's easier online than to sketch in paint.

OpenStudy (anonymous):

\[ -1\le\sin(x)\le1\\ -1\le\sin(\pi x)\le1\\ {1\over4}-1\le{1\over4}+\sin(\pi x)\le{1\over4}+1\\ {-3\over4}\le{1\over4}+\sin(\pi x)\le{5\over4}\\ {-3\over4}\le4^{-x}\le{5\over4}\\ \implies 0\le4^{-x}\le{5\over4}\\ -x\ln(4)\le\ln(5)-\ln(4)\\ (1-x)\le\ln(5) \]

OpenStudy (anonymous):

You will still need to use algebra to get at a lot of values of x that are solutions, but it turns out you can get one just by subbing in.

OpenStudy (amistre64):

im not sure if there is a pretty way to find "all" values that match

OpenStudy (anonymous):

@amistre64 there will be infinite values that "tend" to ge periodic but never are. correct?

OpenStudy (amistre64):

thats pretty much that way i see it yes

OpenStudy (anonymous):

smallest value for "x" is still possible to find though

OpenStudy (anonymous):

hint-> x must be less than \(\pi/2\)

OpenStudy (anonymous):

Well, the answers will approach periodic so you phrase that part of the answer as "for large values, as x -> a + k*pi, sin(x*pi) + 1/4 -> 4^(-x). This should work for all but the first couple of solutions.

OpenStudy (anonymous):

@electrokid do you mean x has to be less than 1/2, since we are looking at sin (x*pi)?

OpenStudy (anonymous):

yes, x = 0178 something